A gardener brought 5 rabbits, after 2 months rabbits became 10, and after 4 months they became 20. If the growth continues on the same ratio, what would be the amount of rabbits after 1 year?
Answer by dkppathak(439)

A gardener brought 5 rabbits, after 2 months rabbits became 10, and after 4 months they became 20. If the growth continues on the same ratio, what would be the amount of rabbits after 1 year?
t1 =5
t2=10
t3=20
series will be
5+ 10 +20 +------ --+13th
this series will be GP
a=5 r= 2 n= 13
general term = ar^n-1
= 5. 2^13-1
=5.2^12
=5x4096 =20480
number of rabbits will be = 20480
total rabbit will be sum of GP =a(r^n -1)/r-1
= 5(2^13 -1)/2-1
= 5(8192-1)/2-1
=5x8191 =40955
answer number of rabbit will be after 1 years =20480 rabbits
total number of rabbit after 1 year =40955 rabbit